# Holographic homogeneous superfluid on the sphere

### v1

@@ 4 Collective modes of holographic superfluid
To make the dispersion relation clearer under spherical homogeneous superfluid, we depict the relationship between real part of modes and angular quantum number $L$ within normal state in Figure 12. We can identify that, considering only the real part of the modes, the dispersion relations in the case of large black hole and small black hole are similar regardless of which of the three channels it is. For the superfluid state, it is presented in Figure 13. From the left column of panels in Figure 13, it can be seen that there both exist a gapless mode for the two types of black holes, which is exactly the hydrodynamic mode. Whether in a superfluid system or normal-fluid system, the transverse gauge field modes are always gapped.
@@
Figure 12: The dispersion relation for normal-fluid state under large black hole (top plots) and small black hole (bottom plots) background. The leftmost plots are from scalar field channel, middle plots are from longitudinal gauge field channel, rightmost plots are from transverse gauge field.
@@
Figure 13: The dispersion relation for superfluid state under large black hole (top plots) and small black hole (bottom plots) background. The left plots are from scalar mode, right plots are from transverse gauge field.
@@
Figure 14: The variation of the real part of low-lying mode with temperature on large black hole background. This mode is exactly the ”first” hydrodynamic excitation of a spherical superfluid according to the reference Topological superfluid .
Last but not least, we observed an interesting dynamic mode with increasing temperature, which is presented in Figure 14. One can observe that this mode decreases initially and then increases as the temperature rises, the turning point coincides precisely with the phase transition point. This collective mode corresponds to the ”first” hydrodynamic excitation within spherical superfluids system Topological superfluid . It probably exists another kind of phase transition, named Berezinskii-Kosterlitz-Thouless(BKT) phase transition Topological superfluid ; BKT , around the superfluid phase transition point.
@@ 5 Summary and discussion
This paper extends the study of holographic superfluids from planar to spherical topology, focusing on phase transitions and collective excitation modes within shell-shaped superfluid systems. The key technique is to expand the gauge field on the sphere using vector spherical harmonics, and then perform linear analysis. This work will lay a foundation for the subsequent study of inhomogeneous superfluid system and turbulence on the shell-shaped system by holographic gravity method.
In this study, we have examined the phase transitions of a holographic homogeneous superfluid system within a shell-shaped geometry, utilizing the Abelian-Higgs model in a Schwarzschild-AdS4 spacetime. The first important discovery is that, for a fixed chemical potential, the critical temperature for the phase transition to the superfluid on the unit sphere is always larger than that for the phase transition to the superfluid on the plane. It still exhibits a normal second-order phase transition for large black hole background, however, the condensation associated with the small black hole as the background geometry demonstrates an anomalous second order phase transition. For the four solutions in the context of two types of black hole backgrounds, we calculated their free energy, respectively. The results indicate that hairy black holes have the greatest thermodynamic stability. In addition, we analyzed the quasi-normal modes in a spherically uniform superfluid at the linear level. The results show that whether in large black hole or small black hole backgrounds, once a superfluid phase transition occurs, unstable modes will always exist in the normal-fluid solution after perturbation. Moreover, although the small black hole is thermodynamically unstable, it is also dynamically stable at least at the linear level. In the study of excitation modes for $l\neq 0$, we find interesting movement of modes associated with temperature and chemical potential. As $l$ varies, this study presents three different channels, which is similar to the case of planar superfluid. Last but not least, at the linear perturbation level, the holographic superfluid system exhibits the ”first” hydrodynamic modes. In this sense, there maybe exist a BKT phase transition process near the superfluid phase transition point BKT , which is an interesting topic to explore. Based on this study, one can get down to discuss the stability and dynamic behavior of inhomogeneous configurations on the spherical superfluid to explore the effects of curved geometry on physical system.
@@ Acknowledgements.
This work is partly supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grant Nos. 12075026 and 12361141825.

### v4

@@ 4 Collective modes of holographic superfluid
To make the dispersion relation clearer under a spherical homogeneous superfluid, we depict the relationship between the real part of frequency and the angular quantum number $L$ for normal fluid state in Figure 12. One can identify that, considering only the real part of the modes, the dispersion relations in the case of large black hole and small black hole are similar regardless of which of the three channels it is. For the superfluid state, it is presented in Figure 13. From the left column of panels in Figure 13, it can be seen that there exist gapless modes for both the two types of black holes, which is exactly the hydrodynamic mode. Whether in a superfluid system or normal fluid system, the modes for transverse gauge field are always gapped. For the gapless mode based on large black hole background, we observed an interesting dynamic behavior with increasing temperature, which is presented in Figure 14. One can observe that this mode decreases initially and then increases as the temperature increases. The turning point coincides precisely with the phase transition point from superfluid to normal fluid with increasing temperature. This collective mode is so called the “first” hydrodynamic excitation within a spherical superfluid system Topologicalsuperfluid .
@@
Figure 12: The dispersion relation in normal fluid state for large (top plots) and small (bottom plots) black holes. The leftmost plots are from scalar field channel, middle plots are from longitudinal gauge field channel, rightmost plots are from transverse gauge field.
@@
Figure 13: The dispersion relation in superfluid state for large (top plots) and small (bottom plots) black holes. The left plots are from scalar mode, right plots are from transverse gauge field.
@@
Figure 14: The variation of the real part of low-lying mode with temperature based on large black hole background. This mode is exactly the “first” hydrodynamic excitation within the spherical superfluid system Topologicalsuperfluid .
@@ 5 Summary and discussion
This paper extends the study of holographic superfluids from planar to spherical topology, focusing on phase transitions and collective excitation modes within the spherical superfluid systems. The key technique is to expand the gauge field on the sphere using vector spherical harmonics and then perform linear analysis. As mentioned in the Introduction part, relevant experimental equipment has already been established in space, and studying Bose-Einstein condensation in a spherical system will be a promising research direction. The work presented in this paper will lay the foundation for using holographic gravity methods to study the spherical Bose-Einstein condensation, opening a window for subsequent research on spherical BECs, such as the previously mentioned issue about Bose-Einstein condensation of interacting bosons in the spherical system. Some of the research findings in this paper can be supported by previous studies BEC-sphere ; Topologicalsuperfluid , i.e. the critical temperature for the phase transition from normal fluid to the superfluid on the unit sphere is always larger than that on the plane, as well as the “first” hydrodynamic excitation mode.
In addition, we identify a normal second-order phase transition with large black hole being the background geometry; however, the condensation associated with the small black hole as the background geometry presents an anomalous second-order phase transition. For the four solutions in the context of two types of black hole backgrounds, we calculated their free energy, respectively. The results indicate that hairy large black hole have the greatest thermodynamic stability. Therefore, in future studies, we will no longer consider small black holes as the geometric background. Moreover, we analyzed the quasi-normal modes in a spherically uniform superfluid at the linear level. Although the small black hole is thermodynamically unstable, it is also dynamically stable at least at the linear level. In the study of excitation modes for $l\neq 0$, we find interesting dynamic behaviors of the modes associated with temperature and chemical potential. As $l$ varies, it presents three different channels, which is similar to the case of the plane.
The phase transition phenomena and dynamical behavior of collective modes discussed in this paper can await experimental verification. This work will build a foundation for the subsequent study of nonlinear structure, such as, soliton, vortex, as well as turbulence, on the unit sphere system with the help of AdS/CFT duality.
@@ Acknowledgements.
This work is partly supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grant Nos. 12075026, 12035016, 12361141825 and 12375058.